OpenStudy (anonymous):

Find the slope of the line whose equation is 4x – 6y = 12 possible answers: -4 -2/3 2/3 6

OpenStudy (skullpatrol):

Have you learned about the "slope-intercept" form of an equation?

OpenStudy (skullpatrol):

\[y =mx +b\]

OpenStudy (owlcoffee):

where "m" represents the slope, and "b" the OY inteception.

OpenStudy (skullpatrol):

@HarukaN are you still there?

OpenStudy (anonymous):

yea i am sorry

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

would that mean 6y=4x + 12? @skullpatrol

OpenStudy (skullpatrol):

You need to put it in the form y = mx + b.

OpenStudy (skullpatrol):

note 1*y = y

OpenStudy (skullpatrol):

$$\Huge 1*y=m*x+b$$

OpenStudy (skullpatrol):

$$\Huge 6*y \neq 1*y$$